Marin Corporation leases a building to Cullumber, Inc. on January 1, 2020. The following facts pertain to the lease agreement.

1. The lease term is 10 years with equal annual rental payments of $3,469 at the end of each year.

2. Ownership does not transfer at the end of the lease term, there is no bargain purchase option, and the asset is not of a specialized nature.

3. The building has a fair value of $34,400, a book value to Marin of $22,000, and a useful life of 15 years.

4. At the end of the lease term, Marin and Cullumber expect the residual value of the building to be $12,400, and this amount is guaranteed by Money, Inc., a third party.

5. Marin wants to earn a 5% return on the lease, and collectibility of the payments is probable.

Prepare the journal entries to record the entries for Marin for 2020 and 2021.

Date Account Titles Debit Credit
1/1/20
12/31/20
12/31/20
Answer #1

Residual value after 10 years =12400

Discounting factor after 10 years @5% =0.6139

Annualised factor for 10 years @ 5% =7.7217

Present value of lease rents = 3469*7.7217 =26,788

Discounted value of residual value = 12400 * .6139 =7612

Total present value = 26787+ 7612 = 34400

No entry as on 1.1.2020

as on 31.12.20

Cash A/c 3469
To Lease Rental Income A/c            3469

(being lease rental income booked for the year)

Depreciation A/c 1467
To asset a/c A/c       1467      

(being depreciation expense booked for the yaer (22000/15))

In the above case, since the lease is operating lease because :

1. not substantial life of asset is given on lease,

2. asset is not transferred at end of lease